At what temperature will 212.7 g of Cl2 expert a pressure of 900 kPa at a volume of 750 mL? Answer in K
1
Expert's answer
2020-03-27T08:58:29-0400

According to the ideal gas law:

p × V = n × R × T

where p - pressure, V - volume, n - number of moles, R -gas constant, T - temperature

From here:

T = p × V / n × R

As n = m/Mr (m - mass, Mr - molecular weight), the final equation is:

T = [p × V × Mr] / [m × R]

As R = 8.3145 J/(mol × K) = 8.3145 (L × kPa)/(mol × K) and Mr(Cl2) = 70.906 g/mol:

T(Cl2) = [900 kPa × 0.750 L × 70.906 g/mol] / [212.7 g × 8.3145 (L × kPa)/(mol × K)] = 27.06 K


Answer: 27.06 K